OpenStudy (anonymous):

You are playing a game in which you flip 3 coins. It costs $1 to play the game, which must be subtracted from your winnings. Calculate the expected value for the game. If all coins show the same (all heads or all tails), you win $9, otherwise, lose your $.

OpenStudy (texaschic101):

$4 bucks per game.....you win 3 games...so that is $12....you subtracted the $3 to play and you have now won $9.

OpenStudy (texaschic101):

you win $4 per game

OpenStudy (texaschic101):

I think :/

OpenStudy (anonymous):

That is correct.

OpenStudy (blockcolder):

Probability of winning is 2/8 (HHH, TTT). You gain $8. Probability of losing is 6/8. You gain -$1. Expected value is (2/8)*(8)+(6/8)*(-1)=2-(6/8)=2/8.